# Squeezing Quantum States in Three-Dimensional Twisted Crystals –ِ

### v1

@@ Acknowledgments
Using these matrix elements, we calculate the energy spectra of particles in the saddle-point potential energy, weakly regularized by the isotropic confining potential. We only include up to a certain maximum value . Physically, this means that there are “hard walls” which prohibit the existence of states beyond a certain radius. In the first calculation, we only include In this case, we find, as shown in Fig. S3, the development of four “edge” states becomes clear as the strength of the saddle-point potential energy increases. If, we include all values up to then the energy spectrum is shown in Fig. S4. There, we also see the “edge” states, but they only become clear for large values of
S5 Boundary Modes of Hofstadter Networks
In the main text, we discussed the bulk spectra of the Hofstadter networks. These networks consist of a periodic array of saddle points described by an intra-saddle-point tunneling amplitude . Tunneling across neighboring saddle points proceeds the inter-saddle-point tunneling amplitude . This is illustrated in Fig. 4 of the main text. In Figs. S5 and S6, we show the spectrum for the bulk and finite-width ribbons of the triangular Hofstadter network, to illustrate the boundary modes appearing in the bulk Hofstadter gaps. We consider both the limit of strong (Fig. S5) and weak (Fig. S6) inter-saddle-point tunneling using the same parameter values as in the main text.

### v2

@@ Appendix D Effect of Saddle-Point Potential Energy
Using these matrix elements, we calculate the energy spectra of particles in the saddle-point potential energy, weakly regularized by the isotropic confining potential. We only include up to a certain maximum value . Physically, this means that there are “hard walls” which prohibit the existence of states beyond a certain radius. In the first calculation, we only include In this case, we find, as shown in Fig. 7, the development of four “edge” states becomes clear as the strength of the saddle-point potential energy increases. If, we include all values up to then the energy spectrum is shown in Fig. 8. There, we also see the “edge” states, but they only become clear for large values of
Appendix E Boundary modes of Hofstadter networks
In the main text, we discussed the bulk spectra of the Hofstadter networks. These networks consist of a periodic array of saddle points described by an intra-saddle-point tunneling amplitude . Tunneling across neighboring saddle points proceeds the inter-saddle-point tunneling amplitude . This is illustrated in Fig. 4 of the main text. In Figs. 9 and 10, we show the spectrum for the bulk and finite-width ribbons of the triangular Hofstadter network, to illustrate the boundary modes appearing in the bulk Hofstadter gaps. We consider both the limit of strong (Fig. 9) and weak (Fig. 10) inter-saddle-point tunneling using the same parameter values as in the main text.